OpenStudy (anonymous):

find the area of the region that lies inside both curves of r=1+cost & r=1-cost

OpenStudy (anonymous):

can you show me the work by integral plz

OpenStudy (anonymous):

its not a line lol its circles

OpenStudy (anonymous):

and i found the limits for t from pi/2 to 3pi/2

OpenStudy (anonymous):

equal both equation: 1+cost=1-cost cost=0

OpenStudy (anonymous):

AREA ?

OpenStudy (anonymous):

I KNO THE T VALUES.. Pi/2 aand 3pi/2

OpenStudy (anonymous):

lol i wasnt talking to you sory :)) and thanks for your help

OpenStudy (anonymous):

for example: \[\Pi/2 \to 3\pi/2\] Then integra it \[\int\limits_{\pi/2}^{3\Pi/2}1+cost -(1-cost)dt\]

OpenStudy (anonymous):

Manuelg your work is not correct, please don't post wrong results delibaretely. That confuses people.

OpenStudy (anonymous):

why is wrong? You are wrong . I'm a theoretical physicist.

OpenStudy (anonymous):

thanks alott ( I wasnt just sure about subtraction or adding btw the areas

OpenStudy (anonymous):

since it says inside both

OpenStudy (anonymous):

dont you get negative answer ???

OpenStudy (anonymous):

the boundary is wrong. You must use : \[-\pi/2 \rightarrow \Pi/2\]

OpenStudy (anonymous):

Area always is positive